# Superadditive Communication with the Green Machine: A Practical Demonstration of Nonlocality without Entanglement .1 Abstract

### v3

@@
Data acquisition The photon passing through the Green Machine is detected by Quantum Opus SNSPD (efficiency 85% at 1550 nm, ideal dark count < 200 per second), tagged by Swabian Instruments TimeTagger Ultra (dead time 2 ns). The detection events of one round are saved to one individual local file along with the synchronization clock ticking at 6.25 MHz. Since the synchronization clock is gated by an extra switch triggered by the computer, the actual communication time is usually less than 50 ms due to the inconsistency of Ethernet latency. During data processing, we manually set the guard band between PPM symbols to be 10 ns for GM1, GM2, and GM3 tests, and 8 ns for GM4 tests. The guarded symbol lengths are 10 ns for GM1, GM2, and GM3 and 2 ns for GM4.
@@
The estimated noise photon rate, including dark counts and ambient light, for GM1, GM2, and GM3 is around per 10 ns guarded symbol over two SNSPDs. For GM4, since its guarded-symbol length is shorter and the dead time after guard-band count affects the register of possible noise photons, we anticipate GM4’s effective noise photon is reduced to per symbol according to our model.
For more details, please see Supplementary Information.
.10 Data availability
All supporting data are available in the manuscript or the supplementary materials. Exemplar raw data and processing code are posted on the online repository Zenodo at https://doi.org/10.5281/zenodo.10215208.
Supplementary Material:
Superadditive Communications with the Green Machine:
A Practical Demonstration of Nonlocality without Entanglement
I I. Historical Green Machine
@@ A Practical Demonstration of Nonlocality without Entanglement
II II. Deep-Space Communications and the Advantageous Range of the Green Machine
To discuss the scenario where our optical Green Machine receiver can show practical advantage, It is necessary to know the received mean photon number and the equivalent distance in current deep-space communication testing platforms. Here, we compare the power budget, i.e., the mean received photon number, of various communication systems deployed in practice and then show the regime where our current GM design can have advantages if being applied in these missions.
A coherent signal with fundamental Hermite-Gaussian mode propagating through a free space link will experience diffraction-induced losses due to the limited aperture size of the receiving system. The power received for aperture limited free space communication link can be compactly written as
with parameters defined in the table I:
@@
For a baseline comparison, we choose to investigate the Lunar Lasercom Demonstration (LLCD) and the Deep Space Optical Communication (DSOC) package operating at Lunar, Martian, and Psyche distances. The relevant parameters for each system are then listed in table II (uplink) and table III (downlink).
@@
Currently, one limiting factor of the pulsed-laser power budget is the load/cost of the customized-high-power Q-switch laser and the gain media it equips. In principle, a customized CW laser could have more average power output, a smaller size, and a lower weight than these pulsed lasers. If we managed to build a GM working at the same modulation rates but using length- BPSK Hadamard codewords encoded over a CW laser beam, the final converted -PPM pulses would attain more photons at higher capacities. We anticipate the GM architecture will be the next paradigm of deep-space communication, especially when the spacecrafts are reaching further, or the human successfully establishes a colony on Mars that requires high-quality, high-capacity, and constant communication. Therefore, it is now necessary to visualize the mean photon for a received PPM pulse for a given communication system as a function of the distance separated. Figure S1 shows the power budget of both LLCD and DSOC platforms’ uplink and downlink and the sweet spot that our current GM4 experiment has shown advantages. For reference, the average distances of the Moon, Mars, Saturn, and Voyager 1 current distance are shown.
III III. Theoretical Backgrounds in Brief
III.1 A. BPSK, PPM, and Hadamard Codewords in Optical Communication
@@
III.2 B. Receivers and Their Capacities
@@
where is the classic mutual information. In our case of optical communication, the transmitter encodes into a light field, and the receiver’s outputs are with photons received per channel per code on average. Then, the PIE is defined as PIE=. Here, output can have one more choice as an erasure channel for no-photon-arriving cases, which is necessary especially when is small.
@@
III.3 C. Superadditive communication and nonlocality without entanglement
@@
Our current experiment has demonstrated superadditive communication and nonlocality without entanglement by showing that a four-stage Green Machine (GM4) receiver decoding length-16 Hadamard BPSK secures a higher PIE than any symbol-by-symbol detections in the photon-starving regime. In the current interpretation, we exclude the net loss of all the receivers. This is a practical distinction for this claim since heterodyne and homodyne receivers usually have much lower losses than the current GM4 setup. However, the current loss of GM4 does not come from any fundamental physical limit, which means it could be improved with better devices, especially the electro-optical modulators. The feasibility of developing low-loss Green Machine receivers will be an important question in forthcoming extensions of this work. Figure S2 shows the PIE simulations of Green Machines with different numbers of stages and other competitors under ideal conditions and with realistic imperfections.
IV IV. Phase Symbol, Channel Phase Noise, and Monte Carlo Simulations
@@
The Monte Carlo results shown in Fig.S3 indicate that the PIEs of Dolinar and homodyne receivers that need an extra phase reference to generate a local oscillator start dropping at 10Hz-level phase noise. Using the same eight-symbol Hadamard encoding and threshold homodyne receiver for decoding will provide more robustness to phase noise but reduce the PIE. On the contrary, the heterodyne receiver can leverage the second quadrature information to correct the phase drift but suffers a reduced PIE. Moreover, if the phase drift is as fast as the modulation bandwidth, the signal arrived at the receiver will be in a mixed state; thus, none of the receivers can extract information from it. To not compromise our claim of practical advantage, we aggressively assume that the upper bound of an ideal heterodyne receiver can retain its PIE until the phase drift approaches the modulation bandwidth with the best possible strategy of phase correction leveraging the extra quadrature. However, in practice, the PIE of the heterodyne receiver will drop at a much slower phase-drift rate due to other practical imperfections.
V V. Experimental Setup and Calibration
Following the Methods section at the end of the main text, here, we expand to the details of the experimental setup (Fig. S4a), the logic flows (Fig. S4b), and the time-multiplexing phase correction techniques (Fig. S5a) in this section. In these figures, we choose to give an example of GM3 since it is at the sweet spot between better visualization and enough complexity for generalization to more stages.
@@
V.1 A. Phase correction of the Green Machine
Let’s take GM3 as an example. During the phase correction of the GM3 (Fig. S5a), AWG1 and AWG2 are turned off. The channel attenuator is also set to zero attenuation so that the full-power CW laser is sent to Bob. Before the GM3 receiver, three concatenated phase modulators (PMs) are used to compensate for the phases of the three stages. The phase correction happens in serial, which means it first starts at S1.
@@
The phase compensation points for all stages can be chosen as a combination of the maximum or minimum of the function, which only changes the code mapping. In Fig. S5a, the detailed flow shows the case where the phase compensation points are all maximum.
We have tested the compensation performance by recording the monitor output with an additional ms delay after ms data acquisition and comparing it with the output during phase scanning. The comparison results of compensating to the maximum at S1, minimum at S2, and maximum at S3 are shown in Fig. S5b. A sampled piece of the histogram is shown in Fig. S5c.
V.2 B. Data acquisition
@@
In the round of one set of the experiment, the channel attenuation is set to the given value (through the serial port command from the computer) after completing phase correction of all stages in the Green Machine. To add channel phase drift, the AWG2 in Fig. S4a will send a corresponding signal to the channel PM. The channel parameter is loaded.
Then, Alice starts sending encoded Hadamard signals to her PM (by AWG1 in Fig. S4a through the serial port command from the computer) to encode the BPSK Hadamard coherent states for 50 ms. After the command is applied, the computer will trigger AWG3 (through the serial port command) to send a high TTL control signal to the switch and let the 6.25 MHz clock signal pass to the Time Tagger. At the same time, the computer will ask Time Tagger to start recording the arrival time for both the SNSPD outputs and the rising edges of the clock signal. Therefore, by compensating for a constant time difference, the clock signal can define which serial of time bins forms one codeword.
After the Time Tagger returns the time-tagged events of all channels, the computer will save the file locally. The histogram of decoding each round of communication is extracted from the corresponding file. From the histogram, the mean photon number received at the SNSPD and the transfer probability matrix for the given codeword are derived accordingly.
V.3 C. Selected and non-selected data
In Fig.2b of the main text, we are showing the non-selected data of multiple trials of all Green Machines, which even includes the data acquired when the Green Machine is occasionally at a bad phase condition. For example, the airflow and the daily activity in the lab sometimes cause unknown movement to the fibers and make the phase unstable. This can be seen from the second and eighth rows of histograms in Fig.2c, where more-than-usual crosstalk is recorded. However, since we have the monitor to check the performance of phase compensation and stability of the interferometers in the Green Machine, we are able to filter out the data acquired when the Green Machine is in a bad phase condition during the ms by confirming it with the monitor outputs (see Fig.S5b). Then, we can discard it and wait for the next round of testing the same codeword.
By doing so, we can extract the data that are not severely affected by the bad phase conditions during continuously running the Green Machine experiments. As shown in Fig.S6a, the crosstalks are reduced to 0.5%, 2.5%, 4.5%, and 7.0% for GM1, GM2, GM3, and GM4 accordingly. Compared to Fig.2b in the main text, this infers how well our Green Machines can work if the long-term phase stability is improved. We also show one example of GM4 at in Fig.S6b to show that low error rates can be obtained for all 16 codewords.

### v4

@@ IV.4 Data acquisition
The photon passing through the Green Machine is detected by Quantum Opus SNSPD (efficiency 85% at 1550 nm, ideal dark count < 200 per second), tagged by Swabian Instruments TimeTagger Ultra (dead time 2 ns). The detection events of one round are saved to one individual local file, along with the synchronization clock ticking at 6.25 MHz. Since the synchronization clock is gated by an extra switch triggered by the computer, the actual communication time is usually less than 50 ms due to the inconsistency of Ethernet latency. During data processing, we manually set the guard band between PPM symbols to be 10 ns for GM1, GM2, and GM3 tests, and 8 ns for GM4 tests. The guarded symbol lengths are 10 ns for GM1, GM2, and GM3 and 2 ns for GM4.
@@
The estimated noise photon rate, including dark counts and ambient light, for GM1, GM2, and GM3 is around per 10 ns guarded symbol over two SNSPDs. For GM4, since its guarded-symbol length is shorter and the dead time after guard-band count affects the register of possible noise photons, we anticipate GM4’s effective noise photon is reduced to per symbol according to our model, resulting in a slower drop of GM4’s PIE at side shown in Fig. 2b. The data processing and the generation of data-related figures in this work are done using Mathematica, Version 14.2,(2024) [Mathematica]
For more details, please see Supplementary Note 5.
V Data availability
The raw detection data of one run of the GM3 that was generated in this study, and the corresponding data processing code have been deposited in the Zenodo available at URL: https://doi.org/10.5281/zenodo.10215208.
Supplementary Information:
Superadditive Communication with the Green Machine as a Practical Demonstration of Nonlocality without Entanglement
X Supplementary Note 1: Historical Green Machine
@@
XI Supplementary Note 2: Deep-Space Communications and the Advantageous Range of the Green Machine
To discuss the scenario where our optical Green Machine receiver can show a practical advantage, it is necessary to know the received mean photon number and the equivalent distance in the current deep-space communication testing platforms. Here, we compare the power budget, i.e., the mean received photon number, of various communication systems deployed in practice and then show the regime where our current GM design can have advantages if applied in these missions.
A coherent signal with a fundamental Hermite-Gaussian mode propagating through a free space link will experience diffraction-induced losses due to the limited aperture size of the receiving system. The power received for aperture-limited free space communication link can be compactly written as
with parameters defined in the Supplementary Table 1:
@@ XI Supplementary Note 2: Deep-Space Communications and the Advantageous Range of the Green Machine
For a baseline comparison, we choose to investigate the Lunar Lasercom Demonstration (LLCD) and the Deep Space Optical Communication (DSOC) package operating at Lunar, Martian, and Psyche distances. The relevant parameters for each system are then listed in Supplementary Table 2 (uplink) and Supplementary Table 3 (downlink).
@@
Currently, one limiting factor of the pulsed-laser power budget is the load/cost of the customized-high-power Q-switch laser and the gain media it equips. In principle, a customized CW laser could have more average power output, a smaller size, and a lower weight than these pulsed lasers. If we managed to build a GM working at the same modulation rates but using length- BPSK Hadamard codewords encoded over a CW laser beam, the final converted -PPM pulses would attain more photons at higher capacities. We anticipate the GM architecture will be the next paradigm of deep-space communication, especially when the spacecrafts are reaching further, or the human successfully establishes a colony on Mars that requires high-quality, high-capacity, and constant communication. Therefore, it is now necessary to visualize the mean photon for a received PPM pulse for a given communication system as a function of the distance separated. Supplementary Figure 1 shows the power budget of both LLCD and DSOC platforms’ uplink and downlink, and the sweet spot that our current GM4 experiment has shown advantages. For reference, the average distances of the Moon, Mars, Saturn, and Voyager 1 current distance are shown.
XII Supplementary Note 3: Theoretical Backgrounds in Brief
XII.1 A. BPSK, PPM, and Hadamard Codewords in Optical Communication
@@ XII Supplementary Note 3: Theoretical Backgrounds in Brief
XII.2 B. Receivers and Their Capacities
@@
where is the classic mutual information. In our case of optical communication, the transmitter encodes into a light field, and the receiver’s outputs are with photons received per channel per code on average. Then, the PIE is defined as PIE=. Here, output can have one more choice as an erasure channel for no-photon-arriving cases, which is necessary, especially when is small.
@@
XII.3 C. Superadditive communication and nonlocality without entanglement
@@
Our current experiment has demonstrated superadditive communication and nonlocality without entanglement by showing that a four-stage Green Machine (GM4) receiver decoding length-16 Hadamard BPSK secures a higher PIE than any symbol-by-symbol detections in the photon-starving regime. In the current interpretation, we exclude the net loss of all the receivers. This is a practical distinction for this claim since heterodyne and homodyne receivers usually have much lower losses than the current GM4 setup. However, the current loss of GM4 does not come from any fundamental physical limit, which means it could be improved with better devices, especially the electro-optical modulators. The feasibility of developing low-loss Green Machine receivers will be an important question in forthcoming extensions of this work. Supplementary Figure 2 shows the PIE simulations of Green Machines with different numbers of stages and other competitors under ideal conditions and with realistic imperfections.
XIII Supplementary Note 4: Phase Symbol, Channel Phase Noise, and Monte Carlo Simulations
@@
The Monte Carlo results shown in Supplementary Figure 3 indicate that the PIEs of Dolinar and homodyne receivers that need an extra phase reference to generate a local oscillator start dropping at 10Hz-level phase noise. Using the same eight-symbol Hadamard encoding and threshold homodyne receiver for decoding will provide more robustness to phase noise but reduce the PIE. On the contrary, the heterodyne receiver can leverage the second quadrature information to correct the phase drift but suffers a reduced PIE. Moreover, if the phase drift is as fast as the modulation bandwidth, the signal arrived at the receiver will be in a mixed state; thus, none of the receivers can extract information from it. To not compromise our claim of practical advantage, we aggressively assume that the upper bound of an ideal heterodyne receiver can retain its PIE until the phase drift approaches the modulation bandwidth with the best possible strategy of phase correction leveraging the extra quadrature. However, in practice, the PIE of the heterodyne receiver will drop at a much slower phase-drift rate due to other practical imperfections.
In a typical free-space laser communications downlink, the phase drift caused by the Earth’s atmospheric turbulence is usually below 100, corresponding to the Greenwood frequency [greenwood1977bandwidth]. The standard adaptive optical system and deformable mirror used in optical telescopes usually have an update rate of 1 kHz (1 frame per ms). Therefore, the maximum Hadamard codeword duration with phase drift immunity is around ms, which corresponds to time bins each ns long (can be converted into PPM of order by a -stage GM). By reducing the bin length with high-bandwidth electronics, the order can be further increased. It is within current technological reach to build a Green Machine for order and be compatible with the state-of-the-art PPM experiments, albeit with K-fold lower peak transmitted power [rielander2023esa, guo2023record].
XIV Supplementary Note 5: Experimental Setup and Calibration
Following the Methods section at the end of the main text, here, we expand to the details of the experimental setup (Supplementary Figure 4a), the logic flows (Supplementary Figure 4b), and the time-multiplexing phase correction techniques (Supplementary Figure 5a) in this section. In these Supplementary Figures, we choose to give an example of GM3 since it is at the sweet spot between better visualization and enough complexity for generalization to more stages.
@@ XIV Supplementary Note 5: Experimental Setup and Calibration
XIV.1 A. Phase correction of the Green Machine
Let’s take GM3 as an example. During the phase correction of the GM3 (Supplementary Figure 5a), AWG1 and AWG2 are turned off. The channel attenuator is also set to zero attenuation so that the full-power CW laser is sent to Bob. Before the GM3 receiver, three concatenated phase modulators (PMs) are used to compensate for the phases of the three stages. The phase correction happens in serial, which means it first starts at S1.
@@
The phase compensation points for all stages can be chosen as a combination of the maximum or minimum of the function, which only changes the code mapping. In Supplementary Figure 5a, the detailed flow shows the case where the phase compensation points are all maximum.
We have tested the compensation performance by recording the monitor output with an additional ms delay after ms data acquisition and comparing it with the output during phase scanning. The comparison results of compensating to the maximum at S1, minimum at S2, and maximum at S3 are shown in Supplementary Figure 5b. A sampled piece of the histogram is shown in Supplementary Figure 5c.
XIV.2 B. Data acquisition
@@
In the round of one set of the experiment, the channel attenuation is set to the given value (through the serial port command from the computer) after completing phase correction of all stages in the Green Machine. To add channel phase drift, the AWG2 in Supplementary Figure 4a will send a corresponding signal to the channel PM. The channel parameter is loaded.
Then, Alice starts sending encoded Hadamard signals to her PM (by AWG1 in Supplementary Figure 4a through the serial port command from the computer) to encode the BPSK Hadamard coherent states for 50 ms. After the command is applied, the computer will trigger AWG3 (through the serial port command) to send a high TTL control signal to the switch and let the 6.25 MHz clock signal pass to the Time Tagger. At the same time, the computer will ask Time Tagger to start recording the arrival time for both the SNSPD outputs and the rising edges of the clock signal. Therefore, by compensating for a constant time difference, the clock signal can define which serial of time bins forms one codeword.
After the Time Tagger returns the time-tagged events of all channels, the computer saves the file locally. The histogram of decoding each round of communication is then extracted from the corresponding file. From the histogram, the mean photon number received at the SNSPD and the transfer probability matrix for the given codeword are derived accordingly.
XIV.3 C. Selected and non-selected data
In Fig.2b of the main text, we show the non-selected data from multiple trials of all Green Machines, which even includes the data acquired when the Green Machine is occasionally in a bad phase condition. For example, the airflow and the daily activity in the lab sometimes cause unknown movement to the fibers and make the phase unstable. This can be seen from the second and eighth rows of histograms in Fig.2c of the main text, where more-than-usual crosstalk is recorded. However, since we have the monitor to check the performance of phase compensation and stability of the interferometers in the Green Machine, we are able to filter out the data acquired when the Green Machine is in a bad phase condition during the ms by confirming it with the monitor outputs (see Supplementary Figure 5b). Then, we can discard it and wait for the next round of testing the same codeword.
By doing so, we can extract the data that is not severely affected by the bad phase conditions while continuously running the Green Machine experiments. As shown in Supplementary Figure 6a, the crosstalks are reduced to 0.5%, 2.5%, 4.5%, and 7.0% for GM1, GM2, GM3, and GM4 accordingly. Compared to Fig.2b in the main text, this infers how well our Green Machines can work if the long-term phase stability is improved. We also include one example of GM4 at in Supplementary Figure 6b to show that low error rates can be obtained for all 16 codewords.
The data processing and the generation of data-related supplementary figures in this work are done using Mathematica, Version 14.2, (2024) [Mathematica].